# Spatio-temporal methods for estimating subsurface ocean thermal response to tropical cyclones

### v2

@@ Abstract
Prevailing theory contends that tropical cyclones (TCs) are primarily driven by wind-induced heat exchange between the air and sea. Accurate characterization of the subsurface ocean thermal response to TC passage is crucial for accurate TC intensity forecasts and climatological modeling, yet that characterization is complicated by the high-noise ocean environment, correlations inherent in spatio-temporal data, relative scarcity of in situ observations, and entanglement of the TC-induced signal with seasonal signals. We present a general methodological framework that addresses these difficulties, integrating existing techniques in seasonal mean field estimation, Gaussian process modeling, and nonparametric regression into a functional ANOVA model. Importantly, we improve upon past work by properly handling seasonality, providing rigorous uncertainty quantification, and treating time as a continuous variable, rather than producing estimates that are binned in time. This functional ANOVA model is estimated using in situ subsurface temperature profiles from the Argo fleet of autonomous floats through a multi-step procedure, which (1) characterizes the upper ocean seasonal shift during the TC season; (2) models the variability in the temperature observations; (3) fits a thin plate spline using the variability estimates to account for heteroskedasticity and correlation between the observations. This spline fit reveals the ocean thermal response to TC passage. Through this framework, we obtain new scientific insights into the interaction between TCs and the ocean on a global scale, including a three-dimensional characterization of the near-surface and subsurface cooling along the TC storm track and the mixing-induced subsurface warming on the track’s right side.
@@ 1 Introduction
Although the methodological framework presented here, in particular the specific technical choices for local estimation of the seasonal mean and ocean variability, is, to our knowledge, novel to the characterization of TC-induced ocean temperature changes, the abstract statistical framework with which we work has deep roots in functional data analysis. In particular, our model for the decomposition of each temperature observation into terms describing the TC-induced signal, a seasonal mean, and ocean variability may be viewed as a functional analysis of variance (functional ANOVA) model (Ramsay and Silverman, 2005), which has previously been used in applications as varied as medicine (Brumback and Rice, 1998; Maldonado et al., 2002), trajectory estimation (Faraway, 1997), and speech (Ramsay et al., 1996). We refer the reader to standard texts for a comprehensive treatment of functional data analysis and spline smoothing in the context of ANOVA decompositions (Ramsay and Silverman, 2005; Gu, 2013).
@@
We contribute to this body of work in the following respects. After first pairing Argo profiles to produce temperature differences relative to the passage of a TC (Cheng, Zhu and Sriver, 2015) [^1] This careful pairing process takes advantage of the highly correlated nature of spatio-temporal data to improve our final temperature change estimates. , we provide methodology to estimate the seasonal shift in upper ocean temperature over the timespan of the profile pairs considered, based upon prior oceanographic work (Ridgway, Dunn and Wilkin, 2002; Roemmich and Gilson, 2009). We find a nontrivial seasonal effect over the timespan of the profile pairs in our analysis, as most TCs occur during months when the subsurface water column experiences warming. Armed with this information, we adjust the temperature differences by removing the estimated seasonal shift. The previous analysis of Cheng, Zhu and Sriver (2015) does not perform this seasonal adjustment, which may impact their estimate of the TC-related signal at larger time lags, as the seasonal warming would not be distinguished from their TC-related signal.
@@
For these seasonally adjusted temperature differences, we further detail a method for estimating their self-variances and cross-covariances, based upon the locally stationary Gaussian process model of Kuusela and Stein (2018). The reasons for fitting these covariances prior to producing the final fits are twofold: ocean dynamics conspire to produce additional heteroskedastic noise at large time lags; and pairs of profiles, especially those nearby in space and time, can be highly correlated. These covariance estimates are then used in a thin plate spline smoother to account for the effect of the correlated, heteroskedastic observations and to provide pointwise confidence intervals. Through these steps, we attain the chief scientific contributions of characterizing the ocean response to TC passage in the continuous time realm, rather than using a binned analysis as in Cheng, Zhu and Sriver (2015), as well as properly accounting for the seasonal variation in ocean temperatures over the time scales of the paired profiles. Moreover, our estimation of within- and between-pair covariances, which were accounted for in (Cheng, Zhu and Sriver, 2015) admits more accurate fits and enables rigorous uncertainty quantification. These statistical procedures are unified in terms of a functional ANOVA model.
@@ 2 Scientific context
A prevailing theory states that tropical cyclones are primarily driven by two physical phenomena: (1) enthalpy flux from the ocean and (2) the difference in temperature between the ocean surface and the tropopause, in a Carnot engine-like process called wind-induced surface heat exchange (WISHE; e.g., Emanuel, 1986, 1999). SSTs of at least $26.5^{\circ}$C are generally sufficient to support TC genesis (McTaggart-Cowan et al., 2015), and additional energy becomes available to the TC as temperatures increase beyond this threshold.
@@
As TCs intensify, their strengthening surface winds induce mixing of the upper ocean, in addition to extracting energy from the ocean surface. This results in a decrease in SST (e.g., a cold wake) and thus a decrease in energy available for further intensification (Bender and Ginis, 2000; Shay, 2010). Therefore, knowledge of subsurface ocean conditions can be important for accurate TC intensity forecasts (e.g., Mainelli et al., 2008). Such information, however, is difficult to obtain through remotely sensed, satellite observations, necessitating the introduction of in situ measurements, such as those provided by the Argo float program (Jayne et al., 2017).
@@
The passage of a TC induces energy transfer from the ocean to the atmosphere via latent heat, inducing temperature decreases at the surface and within the ocean mixed layer (Fisher, 1957; Leipper, 1966; Bender and Ginis, 2000; D’Asaro et al., 2007; Lin et al., 2005; Lloyd and Vecchi, 2011; Balaguru et al., 2012; Elsberry, Fraim and Trapnell, 1976; Price, 1980). Many observational studies, using survey ships or expendable bathythermographs, have examined the thermodynamic response of the ocean surface to TC passage (Price, 1980; Shay and Goni, 2000; Cione and Uhlhorn, 2003; D’Asaro et al., 2007; Dare and McBride, 2011; Haakman et al., 2019), especially before the Argo array of profiling floats was deployed. With more than two million profiles since the early 2000s, Argo floats provide unprecedented spatial and temporal coverage of the global ocean in the upper 2000 decibars (Roemmich and Gilson, 2009; Riser et al., 2016; Jayne et al., 2017). Due to their autonomous nature, they are well-positioned to sample the ocean state in three dimensions before and after the passage of a tropical cyclone. Several previous studies have leveraged Argo profiles to describe TC-related changes in the upper ocean (Liu et al., 2007; Balaguru et al., 2015; Sun et al., 2012; Qu, Song and Maes, ; Cheng, Zhu and Sriver, 2015; Lin et al., 2017; Steffen and Bourassa, 2018; Trenberth et al., 2018), focusing on changes in ocean temperature, salinity, or both, to analyze changes in ocean stratification. The closest work to ours is Cheng, Zhu and Sriver (2015), who perform a binned analysis of time (compared to our continuous treatment of time) and do not account for a seasonal warming effect we discern in our analysis.
@@
Processes regulating the thermal recovery of cold ocean wakes associated with the passage of a TC depend on the parameters of the wake, e.g., depth, width, buoyancy anomaly, and wind stress (Haney et al., 2012). This restratification of the water column following the cold wake results in a net warming of the water column, a warming which increases with TC intensity (Mei et al., 2013). Globally, this net ocean warming is on the order of 0.5 PW (Sriver, Huber and Nusbaumer, 2008; Mei and Pasquero, 2012; Mei and Pasquero, 2013; Mei et al., 2013), i.e., approximately $\nicefrac{{1}}{{6}}$ of the ocean heat transport out of the tropics (Wunsch, 2007). In the long-term mean, this added heat is either advected by ocean currents (e.g., poleward; Emanuel, 2001; Korty, Emanuel and Scott, 2008) or released back to the atmosphere during subsequent cold seasons (Pasquero and Emanuel, 2008; Jansen, Ferrari and Mooring, 2010). For stronger TCs, part of the anomalous warming lies below the winter mixed-layer depth (Mei et al., 2013) and can influence global meridional heat transport. However, the transport of heat may be primarily equatorward (Jansen and Ferrari, 2009). The methodological framework proposed in this manuscript provides an opportunity to contribute to this ongoing discussion, as it helps characterize the evolution of TC-related ocean temperature changes in greater detail, accuracy, and rigor than previously available.
@@ 4.1.2 TC profiles and non-TC profiles
We further partition the Argo profile database into two sets, which separately characterize the ocean state in the presence and absence of TCs. If an Argo profile is within $8^{\circ}$ (latitude and longitude) of a tropical cyclone observation location and within 12-12 days before or +30+30 days after that tropical cyclone observation, we denote it a tropical cyclone profile (TC profile). Any profile not in this set is denoted a non-tropical cyclone profile (non-TC profile). Though the non-TC profiles do not carry any of the TC signal $s(d,\tau)$, they play a crucial role in estimating the seasonal effect $m$ and ocean variability $a$, which characterize the ocean state and dynamics under the important assumption that there are no TC-related effects. These models are fully described in Sections 4.3 and 4.4.
@@ 4.2.1 Pairing process
We follow closely the pairing process introduced in Cheng, Zhu and Sriver (2015). Recall that the tropical cyclone tracks are available at six-hourly time resolution and can be made continuous in time and space through straight line interpolation. Along each of these TC tracks, we seek Argo profiles that reported within $\pm 8^{\circ}$ (longitude/latitude) and within 12-12 days before and +20+20 days after each location on the tropical cyclone track. We partition these profiles into the baseline set (12-12 days before to 2-2 days before) and the signal set (2-2 days before to +20+20 days after). These two sets contain the profiles that are incidental to the TC. The baseline and signal time ranges are chosen primarily to accommodate sampling constraints posed by the Argo network; specifically, most floats sample on a ten-day frequency. Therefore, the baseline time range of 12-12 to 2-2 days ensures that each float can contribute at least one baseline profile. The signal time range allows for at least two signal profiles from a given float (assuming they fit a nearness condition detailed in the sequel). The signal profiles begin at 2-2 days to hedge against storm effects prior to passage and is consistent with the pairing process introduced by Cheng, Zhu and Sriver (2015).
@@ 4.3 Seasonal mean field estimation
to the temperature observations, using all the non-TC profiles that fall within $\pm 8^{\circ}$ (longitude/latitude) of the $(x_{\mathrm{lon}}^{*},x_{\mathrm{lat}}^{*})$ point at which our local regression function is centered.
@@ 4.4 Variability estimates
with $\theta_{\mathrm{lat}},\theta_{\mathrm{lon}},\theta_{t}>0$. As in Kuusela and Stein (2018), these parameters are obtained through the method of maximum likelihood.
@@
$\operatorname{\mathrm{var}}(\tilde{T}_{\text{baseline{}}})=\operatorname{\mathrm{var}}(\tilde{T}_{\text{signal{}}})$ $=\phi(\boldsymbol{x}^{\prime})+\sigma(\boldsymbol{x}^{\prime});$ (15)
@@ 5.4 Depth-cross-track parameterization
For each of these two bins, we average out the temporal component and present the resulting fits as functions of cross-track angle $d$ and pressure level $z$ in Figure 10.
Several noteworthy phenomena are present in Figure 10. In Figure 10, we observe both the primary cooling effect at the ocean and the secondary warming effect in the subsurface due to vertical mixing, as discussed in Sections 5.2 and 5.3. We also observe a sustained cooling effect in Figure 10, during the period called the “recovery” stage in Cheng, Zhu and Sriver (2015). Both results are qualitatively consistent with those in the previous paper.

### v3

@@ Abstract
Tropical cyclones (TCs), driven by heat exchange between the air and sea, pose a substantial risk to many communities around the world. Accurate characterization of the subsurface ocean thermal response to TC passage is crucial for accurate TC intensity forecasts and for an understanding of the role that TCs play in the global climate system. However, that characterization is complicated by the high-noise ocean environment, correlations inherent in spatio-temporal data, relative scarcity of in situ observations, and the entanglement of the TC-induced signal with seasonal signals. We present a general methodological framework that addresses these difficulties, integrating existing techniques in seasonal mean field estimation, Gaussian process modeling, and nonparametric regression into a functional ANOVA model. Importantly, we improve upon past work by properly handling seasonality, providing rigorous uncertainty quantification, and treating time as a continuous variable, rather than producing estimates that are binned in time. This functional ANOVA model is estimated using in situ subsurface temperature profiles from the Argo fleet of autonomous floats through a multi-step procedure, which (1) characterizes the upper ocean seasonal shift during the TC season; (2) models the variability in the temperature observations; (3) fits a thin plate spline using the variability estimates to account for heteroskedasticity and correlation between the observations. This spline fit reveals the ocean thermal response to TC passage. Through this framework, we obtain new scientific insights into the interaction between TCs and the ocean on a global scale, including a three-dimensional characterization of the near-surface and subsurface cooling along the TC storm track and the mixing-induced subsurface warming on the track’s right side.
@@ 1 Introduction
Although the methodological framework presented here, in particular the specific technical choices for local estimation of the seasonal mean and ocean variability, is, to our knowledge, novel to the characterization of TC-induced ocean temperature changes, the statistical framework has close connections to functional data analysis. In particular, for a fixed depth, our model decomposing each temperature observation into terms describing the TC-induced signal, a seasonal mean, and ocean variability may be viewed as a functional analysis of variance (functional ANOVA) model (Ramsay and Silverman, 2005), which has previously been used in applications as varied as medicine (Brumback and Rice, 1998; Maldonado et al., 2002), trajectory estimation (Faraway, 1997), and speech (Ramsay et al., 1996). We refer the reader to standard texts for a comprehensive treatment of functional data analysis and spline smoothing in the context of ANOVA decompositions (Ramsay and Silverman, 2005; Gu, 2013).
@@
The past work closest to our own is Cheng, Zhu and Sriver (2015), which uses an Argo profile pairing procedure to perform a temporally binned analysis of thermal changes in depth and cross-track distance. We build upon this pairing process, which takes advantage of the highly correlated nature of spatio-temporal data, to make the following contributions. First, we provide methodology to estimate the seasonal shift in upper ocean temperature over the timespan of the profile pairs considered, based upon prior oceanographic work (Ridgway, Dunn and Wilkin, 2002; Roemmich and Gilson, 2009). We find a nontrivial seasonal effect over the timespan of the profile pairs in our analysis, as most TCs occur during months when the subsurface water column experiences warming. Armed with this information, we adjust the temperature differences by removing the estimated seasonal shift. The previous analysis of Cheng, Zhu and Sriver (2015) does not perform this seasonal adjustment, which may impact their estimate of the TC-related signal at larger time lags, as the seasonal warming would not be distinguished from their TC-related signal.
@@
For these seasonally adjusted temperature differences, we further detail a method for estimating their self-variances and cross-covariances, based upon the locally stationary Gaussian process model of Kuusela and Stein (2018). The reasons for fitting these covariances prior to producing the final fits are twofold: ocean dynamics conspire to produce additional heteroskedastic noise at large time lags; and pairs of profiles, especially those nearby in space and time, can be highly correlated. These covariance estimates are then used in a thin plate spline smoother to account for the effect of the correlated, heteroskedastic observations and to provide pointwise confidence intervals. Through these steps, we attain the chief scientific contributions of characterizing the ocean response to TC passage in the continuous time realm, rather than using a binned analysis as in Cheng, Zhu and Sriver (2015), as well as properly accounting for the seasonal variation in ocean temperatures over the time scales of the paired profiles. Moreover, our estimation of within- and between-pair covariances, which were not accounted for in Cheng, Zhu and Sriver (2015), admits more accurate fits and enables rigorous uncertainty quantification. These statistical procedures are unified in terms of a functional ANOVA model, which is fit separately over a grid of twenty depth levels.
@@ 2 Scientific context
A prevailing theory states that tropical cyclones are primarily driven by two physical phenomena: (1) energy flux from the ocean and (2) the temperature difference between the ocean surface and the tropopause, the top boundary of the lowest layer of Earth’s atmosphere, in a process called wind-induced surface heat exchange (WISHE; e.g., Emanuel, 1986, 1999). SSTs of at least $26.5^{\circ}$C are generally sufficient to support TC genesis (McTaggart-Cowan et al., 2015), and additional energy becomes available to the TC as SSTs increase beyond this threshold.
@@
As TCs intensify, their strengthening surface winds extract energy from the ocean surface and induce mixing of the upper ocean. This combination of processes decreases SST (creating a cold wake) and thus decreases the energy available for further TC intensification (Bender and Ginis, 2000; Shay, 2010). Also, the magnitude of mixing is proportional to the strength of the TC’s surface winds. Therefore, knowledge of subsurface ocean conditions can impact the accuracy of TC intensity forecasts (e.g., Mainelli et al., 2008). Such information, however, is difficult to obtain through remotely sensed, satellite-based observations, necessitating the introduction of in situ measurements, such as those provided by the Argo float program (Jayne et al., 2017).
Latent heat, the energy needed to evaporate liquid water and released when water vapor condenses, is the primary mechanism by which energy is transferred from the ocean to the atmosphere. Through latent heat transfer, the passage of a TC induces temperature decreases at the ocean surface and within the ocean mixed layer (Fisher, 1957; Leipper, 1966; Bender and Ginis, 2000; D’Asaro et al., 2007; Lin et al., 2005; Lloyd and Vecchi, 2011; Balaguru et al., 2012; Elsberry, Fraim and Trapnell, 1976; Price, 1980). Observational studies have examined the thermodynamic response of the ocean surface to TC passage via observations from survey ships or expendable bathythermographs (Price, 1980; Shay and Goni, 2000; Cione and Uhlhorn, 2003; D’Asaro et al., 2007; Dare and McBride, 2011; Haakman et al., 2019), especially before the Argo array of profiling floats was deployed.
@@
With more than two million profiles since the early 2000s, Argo floats provide unprecedented spatial and temporal coverage of the global ocean in the upper 2000 meters (Roemmich and Gilson, 2009; Riser et al., 2016; Jayne et al., 2017). Due to their autonomous nature, Argo floats are well-positioned to sample the ocean state in three dimensions before and after TC passage. Several previous studies have leveraged Argo profiles to describe TC-related changes in the upper ocean (Liu et al., 2007; Balaguru et al., 2015; Sun et al., 2012; Qu, Song and Maes, ; Cheng, Zhu and Sriver, 2015; Lin et al., 2017; Steffen and Bourassa, 2018; Trenberth et al., 2018), focusing on changes in ocean temperature, salinity, or both, in order to analyze changes in ocean density. The closest work to ours is Cheng, Zhu and Sriver (2015), who perform a binned analysis of temperature changes in time (compared to our continuous treatment of time) and do not account for a seasonal warming effect we discern in our analysis.
@@
Processes regulating the thermal recovery of cold ocean wakes associated with the passage of a TC depend on the parameters of the wake, e.g., depth, width, and wind stress (Haney et al., 2012). Once a TC passes, thermal recovery results in a net warming of the water column, a warming which increases with TC intensity (Mei et al., 2013). Globally, this net ocean warming is on the order of 0.5 PW (Sriver, Huber and Nusbaumer, 2008; Mei and Pasquero, 2012; Mei and Pasquero, 2013; Mei et al., 2013), equivalent to approximately $\nicefrac{{1}}{{6}}$ of the ocean heat transport out of the tropics (Wunsch, 2007). In the long-term mean, this added heat is either advected poleward by ocean currents (Emanuel, 2001; Korty, Emanuel and Scott, 2008) or eventually released into the atmosphere during subsequent cold seasons (Pasquero and Emanuel, 2008; Jansen, Ferrari and Mooring, 2010). For stronger TCs, part of the extra ocean warming occurs below the winter mixed-layer depth (Mei et al., 2013) and can influence global meridional heat transport. However, the transport of this heat may be primarily equatorward (Jansen and Ferrari, 2009), in which case the heat transport would not contribute to the redistribution of heat within the climate system from the warmer tropics to the colder poles. The methodological framework presented in this manuscript provides an opportunity to contribute to this ongoing discussion because it helps characterize the evolution of TC-related ocean temperature changes in greater detail, accuracy, and rigor than previously available.
@@ 4.1.2 TC profiles and non-TC profiles
We further partition the Argo profile database into two sets, which separately characterize the ocean state in the presence and absence of TCs. If an Argo profile is within $8^{\circ}$ (latitude and longitude) of a tropical cyclone observation location and within $12$ days before to $20$ days after that tropical cyclone observation, we denote it a tropical cyclone profile (TC profile). Any profile not in this set is denoted a non-tropical cyclone profile (non-TC profile). Though the non-TC profiles do not carry any of the TC signal $s(d,\tau)$, they play a crucial role in estimating the seasonal effect $m$ and ocean variability $a$, which characterize the ocean state and dynamics under the important assumption that there are no TC-related effects. These models are fully described in Sections 4.3 and 4.4.
@@ 4.2.1 Pairing process
We follow closely the pairing process introduced in Cheng, Zhu and Sriver (2015). Recall that the tropical cyclone tracks are available at six-hourly time resolution and can be made continuous in time and space through straight line interpolation. Along each of these TC tracks, we seek Argo profiles that reported within $\pm 8^{\circ}$ (longitude/latitude) and within $12$ days before to $20$ days after each location on the tropical cyclone track. We partition these profiles into the baseline set (12-12 to 2-2 days relative to the TC) and the signal set (2-2 to +20+20 days relative to the TC). These two sets contain the profiles that are incidental to the TC. The baseline and signal time ranges are chosen primarily to accommodate sampling constraints posed by the Argo network; specifically, most floats sample on a ten-day frequency. Therefore, the baseline time range of 12-12 to 2-2 days ensures that each float can contribute at least one baseline profile. The signal time range allows for at least two signal profiles from a given float (assuming they fit a nearness condition detailed in the sequel). The signal profiles begin at 2-2 days to hedge against storm effects prior to passage and is consistent with the pairing process introduced by Cheng, Zhu and Sriver (2015).
@@ 4.3 Seasonal mean field estimation
to the temperature observations, using all the non-TC profiles that fall within $\pm 8^{\circ}$ (longitude/latitude) of the $(x_{\mathrm{lon}}^{*},x_{\mathrm{lat}}^{*})$ point at which our local regression function is centered. Since these models are fit separately over each window and depth, the estimation procedure can be trivially parallelized.
@@ 4.4 Variability estimates
with $\theta_{\mathrm{lat}},\theta_{\mathrm{lon}},\theta_{t}>0$. As in Kuusela and Stein (2018), these parameters are obtained through the method of maximum likelihood. As with the local mean field, the estimation of these parameters may be done in parallel across windows and pressure levels.
@@
$\operatorname{\mathrm{var}}(\tilde{T}_{\text{baseline{}}})=\operatorname{\mathrm{var}}(\tilde{T}_{\text{signal{}}})$ $=\phi(\boldsymbol{x}^{\prime})+(\sigma(\boldsymbol{x}^{\prime}))^{2};$ (15)
@@ 5.4 Depth-cross-track parameterization
We present the resulting fits as functions of cross-track angle $d$ and pressure level $z$ in Figure 10.
Several noteworthy phenomena are present in Figure 10. In Figure 10, we observe both the primary cooling effect and the secondary warming effect in the subsurface on the right side of the storm due to vertical mixing, as discussed in Sections 5.2 and 5.3. We also observe a sustained cooling effect in Figure 10, during the period called the “recovery” stage in Cheng, Zhu and Sriver (2015). While the primary cooling effect observed in the two time periods is qualitatively consistent with the results presented by Cheng, Zhu and Sriver (2015), their results also show a strong warming effect on both sides of the storm during the recovery stage (Figure 10 in Cheng, Zhu and Sriver, 2015), which is absent from our estimate. Overall, it appears that the recovery stage estimate in Cheng, Zhu and Sriver (2015) is “warm biased” in comparison to our estimate. We suspect that this difference is due to our methodology’s more careful accounting of the seasonal mean shift as described in Section 4.3.